OpenStudy (ujjwal):

show that: (1^2)/2! + (2^2)/3! + (3^2)/4! + (4^2)/5! +...... = e-1

OpenStudy (anonymous):

use taylor series for e^x

OpenStudy (anonymous):

you'll get it

OpenStudy (anonymous):

put x=1

OpenStudy (ujjwal):

@vamgadu i have no idea regarding taylor series

OpenStudy (anonymous):

taylor series expansion of e^x is \[1+x+x ^{2}/2!+x^{3}/3!+....\] now put x=1 in it

OpenStudy (ujjwal):

I don't get the solution that way.. can you try it?

OpenStudy (anonymous):

no no i am wrong waitno i am wrong wait

OpenStudy (anonymous):

is it e-1 or e^-1??

OpenStudy (ujjwal):

It is not e^-1.. its e - 1

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

\[n^{2}/(n+1)!=(1/(n+1)!)-(1/(n)!)+(1/(n-1)!)\]

OpenStudy (anonymous):

use it i think its pretty straight forward now

OpenStudy (experimentx):

try this way!! \[ ((1^2)-1)/2! + ((2^2)-1)/3! + ((3^2)-1)/4! + ((4^2)-1)/5! +...... + 1/2! + 1/3! +1/4! + .. \]

OpenStudy (anonymous):

\[ \frac {n^2} {(n+1)!} = \frac 1{(n-1)!} + \frac 1{(n+1)!}-\frac 1{n!} \] You should be able to do it now.

OpenStudy (anonymous):

note that \[\sum_{1}^{\inf} 1/n! =e-1\]

OpenStudy (ujjwal):

@FoolForMath i get \[n ^{2}/(n+1)! =1/(n-1)! - 1/((n+1)(n-1)!) \] how did you get that result you just posted?

OpenStudy (experimentx):

take lcm!!

OpenStudy (anonymous):

If you do it right, you will get the right answer ;)

OpenStudy (anonymous):

Anyways, if you expand my expression , you will see it's converging to \( e-1\) nicely.

OpenStudy (anonymous):

i dont think you need to expand at all as we are summing each term to infinity, each term give us an e,e-1. and e-2. substitute these in place of summation of reciprocals of (n-1)!,(n)!,(n+1)! respectively. you'll get the answer

OpenStudy (anonymous):

Oh yes that's what I meant. I mean I am too fool to see the e series without expanding few terms :P

OpenStudy (anonymous):

\[\sum_{1}^{\inf}1/(n-1)!=e and \sum_{1}^{\inf}1/(n+1)!=e-2\]

OpenStudy (anonymous):

Yes man I know.

OpenStudy (anonymous):

dont blame your self FFm i know you are a stud. But even the brightest of people make mistakes

OpenStudy (anonymous):

that was meant for @ujjwal

OpenStudy (anonymous):

lol, I am no stud :)

OpenStudy (anonymous):

i know you are

OpenStudy (ujjwal):

thanks everyone!